OpenStudy (anonymous):

After t seconds the position of a particle moving along the x-axis is given (in feet) by x(t) = 5 sin(3t) + 3 cos(2t), t ≥ 0 . Find the acceleration of the particle when t = π.

OpenStudy (tkhunny):

Location: \(x(t)\) -- This is given Velocity: \(\dfrac{d}{dx}x(t) = v(t)\) Acceleration: \(\dfrac{d^{2}}{dx^{2}}x(t) = \dfrac{d}{dx}v(t) = a(t)\) Looks like a couple of derivatives are all that stand in your way!